OpenStudy (anonymous):

Calculate the ROC, dA/dD, where A is the surface area of a sphere of diameter D (the surface area of a sphere of radius r is 4pir^2) I've tried 8pir, and it was incorrect, any ideas?

OpenStudy (turingtest):

r=D/2 A(D)=4(pi)r^2=4(pi)(D/2)^2=(pi)D^2 dA/dD=2(pi)D

OpenStudy (anonymous):

Thank you very much, that worked!